# Real-Time Doppler and Ionospheric Dispersion Correction Techniques for Arbitrary Waveforms Utilizing GPU Compute

### v1

@@ Abstract
General requirements for radar digital signal processing are ionospheric distortion and Doppler dispersion correction, which has historically required radar-specific hardware to implement in real time. Although analog solutions are computationally efficient, they often come with system design drawbacks which limit waveform flexibility and can result in an overall increase of system complexity. With improvements in modern general compute systems, real-time digital signal processing is becoming more realizable using non-radar-specific high-performance compute. In this paper, we present an analysis of general Doppler and ionospheric correction algorithms for arbitrary waveforms for radar digital signal processing. We also include considerations for efficient implementation of these algorithms in software, specifically using GPU hardware. This analysis includes metrics of performance such as execution time and error correction accuracy. We also provide recommendations for application in radar signal processing. We identify two algorithms for dispersion correction: an FFT-based method for ionospheric dispersion and a numerical interpolation method via sinc interpolation for Doppler dispersion. Both of these algorithms are able to compensate for dispersion equivalent in accuracy to waveform-specific analytical methods and were able to be performed in real-time on a single NVIDIA H100 GPU. These methods are waveform agnostic and applied directly to the samples, improving system flexibility and making them easy to incorporate into existing software-defined radio systems.
@@
Daniel J. Vickers is with the Georgia Institute of Technology, Atlanta, GA 30332 USA (e-mail: dvickers6@gatech.edu).
@@ I Introduction
Doppler and ionospheric dispersion generate significant distortions for ground-to-space radar and communication systems. Failure to properly correct for these distortions can result in a decrease in the signal-to-noise ratio (SNR) of a received signal, a loss in timing/location accuracy, and more. Doppler time dilation and ionospheric distortion correction can be solved analytically for specific waveforms [1, 2], but general solutions for arbitrary waveforms are computationally expensive and require flexible compute systems. Historically, systems have been built around the use of a limited selection of radar waveforms, like linear frequency-modulated waveforms (LFMs), to limit the solutions required. Enabling flexible waveforms allows systems to optimize waveforms for various properties, such as range resolution, sidelobe control, Doppler tolerance, etc. [3].This drives the desire to build modern radar systems with the ability to process a wide variety of radar waveforms. Dispersive effects must be compensated for, and this drives the desire for modern radar systems to have a general solution for arbitrary waveform correction, which the paper presents a solution for.
@@
The feasibility of compensating for dispersion is complicated by real-time signal processing requirements. For radar agility, it is desirable to process receive signals on a pulse-to-pulse basis, allowing for dynamic and responsive transmit scheduling and processing [4]. Radar signal processing has historically been implemented in radar-specific hardware such as field-programmable gate arrays (FPGAs), analog components, and application-specific integrated circuits (ASICs) due to their high signal processing rates [5]. Although radar-specific hardware provides optimal performance, it often requires longer development times, higher costs, and less flexible system designs. Software flexibility improvements can be obtained using more general computing systems such as central processing units (CPUs) and graphics processing units (GPUs), which could allow for more sophisticated digital signal processing (DSP) solutions [6]. The growth trends of computational power for both CPUs and GPUs have been observed [7]. As on-chip processing power has improved, software-defined radios (SDRs) that utilize general compute systems have become more common in research, communication, astronomy, defense, and hobbyist use [8, 9, 10].
@@
In this paper, we address the issue of real-time radar signal processing on arbitrary waveforms by utilizing general GPU compute to compensate for dispersive effects. We do so by presenting correction methods of ionospheric and Doppler dispersion for SDR. We also explore computational optimizations of these algorithms on a GPU, enabling real-time processing. Documentation of these optimizations and their performance on an NVIDIA H100 GPU are presented here.
We review analytic methods for dispersion compensation of LFMs in sections I-A and I-B. We then present methods of arbitrary ionospheric and Doppler dispersion correction in section II. In order to optimize these algorithms for system flexibility, the methods of arbitrary ionospheric and Doppler dispersion correction presented are numerical methods that can be applied to any discretely-measured signal with no information of the modulation type. We will then compare the arbitrary dispersion correction methods to analytic methods for LFMs in section III, and conclude with a discussion of our study in section IV. A list of all variables used throughout the paper can be found in Table I.
TABLE I: A list of all variables used throughout this paper.
Variable Meaning
$q_{e}$ charge of an electron
$m_{e}$ mass of an electron
$\epsilon_{0}$ permittivity of free space
$c$ speed of light
$f$ frequency
$f_{0}$ waveform start frequency
$E$ total electron content (TEC) along the path of propagation
$K_{2}$ plasma delay constant, linearly proportional to $E$
$B$ waveform bandwidth
$T$ waveform pulse width
NN number of received waveform samples
$v_{r}$ range rate of target
$t$ time
I-A Ionospheric Dispersion Compensation for LFMs
The ionosphere is a relatively plasma-dense layer of the upper atmosphere, starting around an altitude of approximately 50 km. The density of the plasma in this atmospheric layer can fluctuate, but is typically densest in increased sunlight, such as during the day, during the summer, and along the equator[11]. As an electromagnetic signal propagates through the ionosphere, the plasma causes a dispersive effect on the signal. A review of ionospheric effects on radio signals can be found in Appendix A. In brief, ionospheric distortion of a waveform is modeled as a frequency-based time delay of the signal, given as
$\begin{split}\tau(f)=\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}E=\frac{K_{2}}{cf^{2}}.\end{split}$ (1)
In (1), $E$ is the total electron content (TEC) along the path of propagation and $K_{2}$ is the plasma delay constant, $q_{e}$ and $m_{e}$ are the charge and mass of an electron, $\epsilon_{0}$ is the permittivity of free space, $c$ is the speed of light, and $f$ is the frequency. The TEC increases when the plasma density of the ionosphere is higher and when the signal propagates through more of the ionosphere. Equation 1 is computed via an approximation that should remain valid for any radar band at HF frequencies or higher. Because the time delay is inversely related to the frequency of the waveform, ionospheric dispersion is particularly impactful on SNR loss and range error at lower-frequency radar bands, such as HF, VHF, UHF, and L. The SNR loss due to dispersion modeled from (1) is also particularly significant for wide-bandwidth signals. Thus, these effects can remain relevant at higher frequency bands given a waveform with sufficient bandwidth.
@@ I-A Ionospheric Dispersion Compensation for LFMs
Ionospheric dispersion correction for waveforms has been performed previously with various methods. High-accuracy methods for correction approximate solutions calculated for specific waveforms [1]. This analytic method can be ideal for cases where known simple waveforms will be transmitted. However, the choice to use an analytic method typically reduces system flexibility. For numerical approximations, another class of methods involves using finite impulse response (FIR) filters with a frequency-dependent response matched to the ionosphere [12, 13, 14]. An overview of modern FIR filtering methods for ionospheric dispersion correction can be found in Ref. [15]. These FIR filters typically operate by representing the ionospheric distortion as a frequency response with three unique input coefficients. All but one of the coefficients are held constant while the variable coefficient is tuned to optimal performance. These methods are computationally efficient, but are not exact. FIR filters are commonly best for applications where one or more parameters of the frequency response are unknown.
@@
To correct for the ionospheric distortion of a waveform, a frequency-dependent correction can be applied by temporally translating the waveform $2\tau(f)$ (to account for the two-way propagation). For continuous waveforms (CWs) with constant frequency, ionospheric effects are reduced to a uniform translation in time (or phase). This problem becomes nontrivial when correcting for waveforms with frequency modulation and more complex frequency characteristics.
@@
A known method to analytically compensate for ionospheric dispersion of LFMs is to apply a correction via predistortion, where the transmitted waveform is inversely distorted in relation to the predicted ionospheric effect. The goal of predistortion is to compensate for the ionosphere on transmit such that the desired waveform is received. This method requires analytic solutions or approximations for any given waveform type. For LFMs, the desired frequency of the transmitted waveform is given as
@@
In (2) $T$, $B$, and $f_{0}$ are the pulse width, bandwidth, and start frequency of the LFM, respectively. One method of solving (2) was presented by Halpin, Urkowitz, and Maron [1]. Their result approximates the frequency as a polynomial expansion of the form
@@
There are many ways to determine the coefficients of (3). Here we will note the Pseudo–Chebyshev polynomial expansion method [^1] Note that the results presented in (7) and (8) differ from those presented in Ref. [1] by a factor of 2 on two terms. We concluded that this factor of 2 difference is a minor error and elaborate on this result in Appendix B. from Ref. [1] in which the coefficients $\mu_{0}$, $\Delta\mu$, and $\gamma$ are calculated as
@@
Equation (9) cannot be integrated directly to compute the phase of predistorted waveform. However, the phase can be integrated numerically to provide another approximation method for correction of ionospheric effects via predistortion with error related to the method of numerical integration. Both of the predistortion methods shown in (3) and (9) provide useful points of comparison for accuracy of ionospheric correction methods for LFMs.
I-B Background of Doppler Dispersion Correction
Doppler dispersion of a waveform is exactly modeled by the two-way longitudinal Doppler shift where the signal is time dilated such that
$S_{doppler}(t)=S(\alpha t).$ (13)
Here $\alpha=\frac{1+v_{r}/c}{1-v_{r}/c}$, and $v_{r}$ is the range rate of the target [16]. Note that we use the convention where $v_{r}$ is positive when the target is approaching the sensor. For a given signal, $S(t)$, the distorted waveform is given as:
The SNR loss and range error due to (13) is primarily a function of bandwidth and modulation type. As the bandwidth of a waveform increases, so does its Doppler-dispersion error. LFMs are considered relatively Doppler tolerant, but other modulation types will typically incur more error.
For narrow-band signals, it can be sufficient to approximate (13) via frequency conversion (also known as time domain downconversion). In frequency conversion, the signal is multiplied by a tone in the time domain, resulting in an upconversion or downconversion. This algorithm is extremely computationally efficient as it scales with a time complexity [^2] For this paper, we present all computational time complexity and processing error in “Big-O” ($\mathcal{O}$) notation. This notation is common in the fields of physics and computational science, but less so in radar DSP. In this notation, the lead-order complexity of a polynomial describing some performance metric is given. For time complexity, lower order compute times are desirable. For example, a time complexity of $\mathcal{O}(N)$ is considered more optimal than $\mathcal{O}(N^{2})$. of $\mathcal{O}\left(N\right)$. However, the frequency shift is only exact at one frequency, and induces errors for pulses with large bandwidths and/or steep phase slopes [17]. For LFMs, this error scales as $\mathcal{O}\left(B^{2}\right)$, where $B$ is the bandwidth of the signal. For arbitrary waveforms, one typically expects this error to be more impactful.
For any signals with a known analytic representation, (13) can be calculated exactly by resampling the signal according to the target range rate. More specifically, when correcting LFMs, this can be modeled by generating a new LFM with a modified phase slope and pulse width according to (13). This analytic resampling method is typically ideal because of its exact correction and because the time complexity scales as $\mathcal{O}\left(N\right)$. However, when performing DSP, the transmit and receive signals are discretized by sampling the waveform. It can then become desirable to pass these samples between software interfaces in order to simplify the overall software complexity of the system. In these cases where no analytic representation of the signal is available, one is forced to use the general solution to approximate the resampling numerically. There are many known ways to compute this numerical resampling. For this paper, we consider the following methods:
  • Polynomial Interpolation: A linear, cubic, higher order, or spline interpolation function is used. While this method is computationally efficient, typically scaling as $\mathcal{O}\left(N\right)$, it results in a loss of power in the interpolated signal due to interpolation error. Higher sample rate signals reduce this error, but raise the computational load.
  • FFT P/Q Resampling: The Fourier transform of the signal has terms removed or added followed by an inverse Fourier transform is taken. This results in the exact resampling assuming a Nyquist-limited waveform. This method can be computationally expensive due to requiring discrete Fourier transforms (DFTs) of non-ideal numbers of samples. This leads to variable sample rates that depend upon the specific number of sample used, and scales as $\mathcal{O}\left(N\log{N}\right)$ in the ideal case. This method is also only exact to an integer number of samples added or removed. A brief exploration of the accuracy and performance of FFT resampling can be found in Ref. [18].
  • Sinc Interpolation: Allows exact resampling of Nyquist limited signals. The exact form of sinc interpolation grows as $\mathcal{O}\left(N^{2}\right)$ in time, but windowing the filter to a size of $N_{window}$ samples results in a time complexity of $\mathcal{O}\left(N\times N_{window}\right)$. The choice of features such as taper window size, taper, and initial sampling rate determine the performance of this algorithm. An overview of Whittaker–Shannon interpolation and general interpolation of Nyquist-limited signals can be found in Ref. [19].
II Methods
We sought a general solution to previous dispersion correction methods by leveraging modern hardware to enable algorithms that were previously too computationally intense. In this section, we present arbitrary methods of Doppler and ionospheric dispersion correction. In Section II-A, we present a numeric method for correcting ionospheric dispersion via a strict time delay from (1) which is applied in the frequency domain. In Section II-B we use (13) to represent Doppler dispersion as an interpolation and explore various methods of interpolation valid for GPU compute.
II-A Ionospheric Dispersion Compensation Via FFT Phase Correction
@@
In this section, we propose a method for numerically determining the ionospheric distortion correction of a transmitted or received waveform in real time directly from discrete samples. Recall from the time-shifting property of the Fourier transform that any time shift in the time domain is represented as a phase shift in the frequency domain. One can apply a frequency-dependent phase shift for each frequency bin using the temporal translation from (1). Thus, using a time delay of $2\tau(f)$ (to account for two-way ionospheric dispersion) one can conclude that the ionospheric effect on a transmitted waveform, $S_{Tx}$, due to two-way propagation will result in a received waveform, $S_{Rx}$, of
$\begin{split}S_{Rx}&=\mathcal{F}^{-1}\left[\mathcal{F}(S_{Tx})\exp\left(2\pi i(2\tau)f\right)\right]\\ &=\mathcal{F}^{-1}\left[\mathcal{F}(S_{Tx})\exp\left(\frac{4\pi iK_{2}}{cf}\right)\right].\end{split}$ (14)
@@
On its own, (14) is a convenient algorithm to directly apply ionospheric distortion to arbitrary waveforms in software. Equation (14) can be applied to the transmitted signal to use as the matched filter during pulse compression. Solving for the inverse of (14) results in a method to remove the ionospheric distortion accrued during propagation and is identical to applying a temporal shift of $2\tau(f)$. The inversion of (14) yields
$\begin{split}S_{Tx}=\mathcal{F}^{-1}\left[\mathcal{F}(S_{Rx})\exp\left(\frac{-4\pi iK_{2}}{cf}\right)\right].\end{split}$ (15)
@@
A similar method to (15) was applied by Peters, Schroeder, and Romero-Wolf [20] for post-processing of satellite data when determining the TEC values of Europa’s atmosphere. Here (15) is intended to correct for the ionosphere in real time given a known TEC value. The error of this algorithm is linearly related to the frequency resolution in the frequency domain. By the Nyquist-Shannon sampling theorem, this means that the error scales as $\mathcal{O}(t_{range})$, where $t_{range}$ is the total length of time of the range window being processed.
@@
In practice, the Fourier transformations used in (15) must be done discretely as Fast Fourier Transforms (FFTs) and inverse FFTs (iFFTs). In the past, it has been too slow to perform the FFT and iFFT in real-time applications on high-bandwidth signals. Modern improvements in hardware and algorithms have made this computation achievable to implement digitally in real time.
@@
Compared to using a polynomial expansion technique, this proposed method offers several benefits. The technique shown in (15) is a direct calculation of the frequency dependent phase shift. This results in a few lines of code, allowing a first step in prototyping with little chance for errors. This technique can distort a waveform on transmit (predistortion) or on receive (post-distortion), and is valid for arbitrary waveforms. The algorithm’s only requirement is that the waveform is sampled within the Nyquist limit. When an analytical solution is inaccessible, (15) can be used as a general numerical method.
@@
The primary drawback of (15) is that it is more computationally expensive in cases where an analytic method can be used. This high computational cost is primarily due to the use of the FFT and iFFT algorithms. The time complexity of (15) scales as $\mathcal{O}\left(N\log{N}\right)$ when the number of samples is exactly a power of two ($N=2^{n}\text{ for }n\in\mathbb{Z}$) and scales as much as $\mathcal{O}\left(N^{2}\right)$ in other cases. The only way to mitigate this cost is with more processing power. The increased processing requirement is why (15) is an excellent candidate for parallelization on a GPU. The computational work done in the frequency domain can also be bundled with other calculations that require FFTs in the signal processing chain and benefit from the efficiency of the shared use of FFT compute time. An example flow diagram of how one could combine pulse compression with (15) for post-distortion is shown in Fig. 1. These optimizations can make our method viable in SDRs to be used for real-time processing in radars and communication systems.
@@
II-B Numeric Doppler Dispersion Correction Via Interpolation

### v2

# Real-Time Doppler and Ionospheric Distortion Correction Techniques for Arbitrary Waveforms Utilizing GPU Compute

@@ Abstract
A general requirement for digital signal processing is ionospheric distortion and Doppler dispersion correction, which has historically required radar-specific hardware to implement in real time. With improvements in modern general compute systems, real time digital signal processing is becoming more realizable using non-radar-specific high performance compute. In this paper, we present an analysis of general Doppler correction algorithms for arbitrary waveforms and a novel ionospheric correction algorithm for radar digital signal processing. We also include considerations for efficient implementation of these algorithms in software, specifically using GPU hardware.
@@
Daniel J. Vickers is with MIT Lincoln Laboratory, Lexington, MA 02421 USA (e-mail: Daniel.Vickers@ll.mit.edu).
@@ I Introduction
Doppler and ionospheric dispersion generate significant distortions for ground-space radar and communication systems. Failure to properly correct for these distortions can result in a decrease in the signal-to-noise ratio (SNR) of a received signal, a loss in timing/location accuracy, and more. Doppler time dilation and ionospheric distortion correction can be solved analytically for specific waveforms [1], but general solutions for arbitrary waveforms are computationally expensive and require flexible compute systems. Radar signal processing has historically been implemented in radar-specific hardware such as field-programmable gate arrays (FPGAs), analog components, and application-specific integrated circuits (ASICs) due to real-time processing requirements. While radar-specific hardware provides optimal performance, it often requires longer development times, higher costs, and less-flexible system designs. Software flexibility improvements can be obtained using more-general compute systems like central processing units (CPUs) and graphics processing units (GPUs), which could allow for more-sphistocated digital signal processing (DSP) solutions. The growth trends of computational power for both CPUs and GPUs has been observed [2]. As on-chip processing power has improved, software-defined radios (SDRs) leveraging general compute systems have become more common in research, communication, astronomy, defense, and hobbyist use [3, 4, 5].
@@
In this paper, we present ionosphere and Doppler dispersion correction methods for use in radar DSP to be performed using non-radar-specific compute for arbitrary waveforms. We also also explore computational optimization of these algorithms for SDR on a GPU for use in real-time processing. Documentation of optimizations and their performance on an NVIDIA H100 GPU are presented here.
Doppler dispersion of a waveform is exactly modeled by the two-way longitudinal relativistic Doppler shift where the frequency, $f$, is dilated such that
$f_{doppler}=\frac{f}{\alpha}.$ (1)
Here $\alpha=\frac{1+v_{r}/c}{1-v_{r}/c}$, $v_{r}$ is the range rate of the target, and $c$ is the speed of light [6]. Note here that we use the convention that $v_{r}$ is positive when the target is approaching the sensor. For a given signal, $S(t)$, the doppler dispersion distorted waveform is given as:
$S_{doppler}(t)=S(\alpha t).$ (2)
For any signals with a known analytic representation, (2) can be computed exactly by resampling the signal according to the target range rate. More specifically, when correcting linear frequency modulated waveforms (LFMs), this can be modeled by generating a new LFM with a modified phase slope according to (1). This analytic resampling method is typically ideal because of its exact correction and because the time complexity scales as $\mathcal{O}\left(N\right)$. However, when performing DSP, the transmit and receive signals are discretized by sampling the waveform. It can then become desirable to pass these samples between software interfaces in order to simplify the overall software complexity of the system. In these cases where no analytic representation of the signal is available, one is forced to use the general solution of approximating the resampling numerically. There are many known ways to compute this numerical resampling. For this paper, we considered these methods:
  • Time Domain Downconversion or Frequency Conversion: The signal is multiplied by a tone in the time domain, resulting in an upconversion or downconversion. This algorithm is extremely computationally efficient as it scales with a time complexity of $\mathcal{O}\left(N\right)$. However, the frequency shift is only exact at one frequency, and induces errors for pulses with large bandwidths and/or steep phase slopes [7]. For LFMs, this error scales as $\mathcal{O}\left(B^{2}\right)$, where $B$ is the bandwidth of the signal. For arbitrary waveforms, one typically expects this error to be more-severe.
  • Polynomial Interpolation: A linear, cubic, higher order, or spline interpolation function is used. While this method is computationally efficient, typically scaling as $\mathcal{O}\left(N\right)$, it results in a loss of power in the signal, especially at turnovers in the value. Higher sample rate signals can result in less error, but raise the computational load.
  • FFT P/Q Resampling: The Fourier transform of the signal has terms removed or padded and then the inverse Fourier transform is taken. This results in the exact resampling assuming a Nyquist limited waveform; however, this method can be computationally expensive due to requiring discrete Fourier transforms (DFTs) of non-ideal numbers of samples. This leads to variable sample rates that depend upon the specific number of sample used, and scales as $\mathcal{O}\left(N\log{N}\right)$ in the ideal case. This method is also only exact out to a integer number of samples to pad or remove. A brief exploration of the accuracy and performance of FFT resampling can be found in Ref. [8].
  • Sinc Interpolation: Allows exact resampling of Nyquist limited signals. The exact form of sinc interpolation grows as $\mathcal{O}\left(N^{2}\right)$ in time, but windowing the filter to a size of $N_{window}$ samples results in a time complexity of $\mathcal{O}\left(N\times N_{window}\right)$. The choice of taper window size, taper, initial sampling rate, and other features determines the performance of this algorithm. An overview of Whittaker-Shannon interpolation and general interpolation of Nyquist limited signals can be found in Ref. [9].
@@
Ionosphere distortion correction for waveforms has been performed previously with various methods. One high-accuracy method approximates solutions calculated for specific waveforms [1]. For the same reason as Doppler dispersion correction, this analytic method can be ideal for cases where known simple waveforms will be transmitted. However, the choice to use an analytic method typically increases system complexity and reduces system flexibility. For numerical approximations, another class of methods involves using finite impulse response (FIR) filters with a frequency dependent response matched to the ionosphere [10, 11, 12]. An overview of modern FIR filtering methods can be found in Ref. [13]. These FIR filters typically operate by representing the ionospheric distortion as a frequency response with three unique input coefficients. All but one of the coefficients are held constant while the variable coefficient is tuned to optimal performance. These methods are computationally efficient, but are not exact and are commonly best for applications where one or more parameters of the frequency response are unknown.
In this paper, we will present an ionosphere distortion correction method that is performed computationally and is valid for arbitrary waveforms. We will also present an analysis of various Doppler dispersion correction methods via interpolation for arbitrary waveforms. We will be leveraging modern hardware to explore the implementation of these algorithms in real time on a GPU as well as provide an documentation of various optimization strategies considered for implementation of these algorithms.
II Ionospheric Distortion Correction
A review of ionospheric effects on radio signals can be found in Appendix A. In brief, ionospheric distortion of a waveform is modeled as a frequency-based time delay of the signal, given as
$\begin{split}\tau(f)=\frac{e^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}E=\frac{K_{2}}{cf^{2}}.\end{split}$ (3)
In (3), $E$ is the total electron content (TEC) along the path of propagation and $K_{2}$ is the plasma delay constant.
@@ II Ionospheric Distortion Correction
To correct for the ionospheric distortion of a waveform, an inverse time shift can be applied by temporally translating the waveform $2\tau(f)$ (to account for the two-way propagation). For continuous waveforms (CWs) with constant frequency, ionospheric effects are reduced to a uniform translation in time (or phase). This problem becomes non-trivial when correcting for waveforms with frequency modulation and more complex frequency characteristics. To perform real-time correction of ionospheric effects for receive waveforms, we sought a simplified approach to previously used methods. By leveraging modern hardware to perform more computationally intense algorithms, our goal was a more-general correction method that results in an overall simpler algorithm and software interface. In this section, we present a mathematic background of the ionospheric correction methods. We then present our numeric method for correction ionospheric distortion and compare this numeric method to analytic methods. Finally, we discuss the time performance of this method.
II-A Ionospheric Dispersion Compensation Via Predistortion
@@
A known method to analytically compensate for ionospheric dispersion is to apply a correction via predistortion, where the transmitted waveform is inversely distorted in relation to the predicted ionospheric effect. The goal being to compensate for the ionosphere such that the desired waveform is received. This method requires precise analytic solutions or approximations for any given frequency-modulated waveform. One of these predistortion methods was developed for LFMs, and was presented by Halpin, Urkowitz, and Maron [1]. The derivation of their polynomial approximation can be found in Appendices A and C of Ref. [1]. Their result is repeated here in (4):
@@
Here $T$, $B$, and $f_{0}$ are the pulse width, bandwidth, and start frequency of the LFM, respectively. (4) has typically been solved by approximating the frequency as a polynomial expansion of the form
@@
There are many ways to determine the coefficients of (5). Here we will note the Pseudo-Chebyshev polynomial expansion method [^1] Note that the results presented in (9) and (10) differ from those presented in Ref. [1] by a factor of 2 on two terms. We concluded that this factor of 2 difference is a minor error and elaborate on this result in Appendix B. from Ref. [1] in which the coefficients $\mu_{0}$, $\Delta\mu$, and $\gamma$ are calculated as
@@
Recall that we can represent a phase modulated waveform as $s(t)=\exp[i\phi(t)]$, where $\phi(t)=2\pi\int_{0}^{t}f(t^{\prime})dt^{\prime}$. The case of (11) is too complicated to integrate directly and yield an analytic representation of the predistorted waveform’s phase. However, the phase can be integrated numerically to provide another approximation method for correction of ionospheric effects via predistortion with error related to the method of numerical integration. Both of the predistortion methods shown in (5) and (11) provide useful points of comparison for accuracy of ionospheric correction methods.
II-B Ionospheric Dispersion Compensation Via FFT Phase Correction
@@
In this section, we propose a method for numerically determining the ionospheric distortion correction of a transmitted or received waveform in real time directly from discrete samples. Recall that given a function $g(t)$, we represent the Fourier transform as $\hat{g}(f)=\int{g(t)\exp(-2\pi itf)dt}=\mathcal{F}(g(t))$. Note that for any function, if $g(t-t_{0})=h(t)$ then $\hat{g}(f)\exp(-2\pi it_{0}f)=\hat{h}(f)$. If $t_{0}$ is a constant, then $\exp(-2\pi it_{0}f)$ is a simple tone, and $h(t)$ is a linear translation by time $t_{0}$. If instead, we make $t_{0}$ a function of frequency, we can translate each frequency in our function by a frequency dependent amount. Using the temporal translation from (3) we can conclude that the ionospheric effect on a transmitted waveform, $s_{trans}$, due to two-way propagation will result in a received waveform, $s_{rec}$, of
$\begin{split}s_{rec}&=\mathcal{F}^{-1}\left[\mathcal{F}(s_{trans})\exp\left(2\pi i(2\tau)f\right)\right]\\ &=\mathcal{F}^{-1}\left[\mathcal{F}(s_{trans})\exp\left(\frac{4\pi iK_{2}}{cf}\right)\right].\end{split}$ (15)
@@
On its own, (15) is a convenient algorithm to directly apply ionospheric distortion to arbitrary waveforms in software. Solving for the inverse of (15) results in a method to remove the ionospheric distortion accrued during propagation and is identical to applying a temporal shift of $2\tau(f)$. The inversion of (15) yields
$\begin{split}s_{trans}&=\mathcal{F}^{-1}\left[\mathcal{F}(s_{rec})\exp\left(2\pi i(-2\tau)f\right)\right]\\ &=\mathcal{F}^{-1}\left[\mathcal{F}(s_{rec})\exp\left(\frac{-4\pi iK_{2}}{cf}\right)\right].\end{split}$ (16)
@@
A similar method to (16) was applied by Peters, Schroeder, and Romero-Wolf [14] for post-processing of satellite data when determining the TEC values of Europa’s atmosphere. Here, we intend for the algorithm to correct for the ionosphere in real time given a known TEC value.
@@
In practice, the Fourier transformations used in (16) must be done discretely as Fast Fourier Transforms (FFTs) and inverse FFTs (iFFTs). Performing the FFT and iFFT for each received waveform has been too slow to solve in the past for real time applications on high-bandwidth signals. Modern improvements to hardware and algorithms have made this computation achievable to implement digitally in real time for high bandwidth signals.
@@
Compared to using a polynomial expansion technique, our proposed method offers several benefits. The technique shown in (16) is a direct calculation of the frequency dependent phase shift. This results in a few lines of code, allowing a first step in prototyping with little chance for errors. Our technique can distort a waveform on transmit (predistortion) or on receive (post-distortion), and is valid for arbitrary waveforms. The algorithm’s only requirement is that the waveform is sampled within the Nyquist limit. When an analytical solution is inaccessible, (16) can be used as a general numerical method.
@@
The primary drawback of our proposed method is that it is more computationally expensive in cases where a polynomial expansion method can be used. This high computational cost is primarily due to the use of the FFT and iFFT algorithms. This cost can be mitigated with the use of modern hardware. The algorithm shown in (16) is also an excellent candidate for parallelization on a GPU. The computational work done in the frequency domain can also be bundled with other calculations that require FFTs in the signal processing chain and benefit from the efficiency of the shared use of FFT compute time. An example flow diagram of how one could combine pulse compression with (16) for post-distortion is shown in Fig. 1 These optimizations can make our method viable in SDRs to be used for real-time processing in radars and communication systems.
@@
II-C Comparison to Ionospheric Correction Predistortion Methods
We tested our method against results for LFM predistortion using the Pseudo-Chebyshev approximation method and the cubic solution from (11) using trapezoidal integration to determine the phase. We chose the Pseudo-Chebyshev method because it generally has one of the highest SNRs for LFM predistortion methods. Results from one of our sample test cases can be seen in Fig. 2.
Figure 2: Comparison of our FFT-based method with both the undistorted waveform (labeled LFM), the Pseudo-Chebyshev polynomial approximation (labeled POLY), our FFT based method (labeled FFT), and our exact frequency from the cubic formula shown in Eq. 11 integrated via trapezoidal integration (labeled CUBIC). This test was run with $f_{0}=413~\text{MHz}$, $B=18~\text{MHz}$, a sample rate of $f_{s}=2.048~\text{GHz}$, $E=100e16~\text{electrons}/\text{m}^{2}$, and $T=100~\mu\text{s}$. As we can see, compared to the total amount of energy in our original signal, shown in the $\text{LFM}\circledast\text{LFM}$ curve, we lose little SNR for the $\text{CUBIC}\circledast\text{POLY}$ curve, which is the cross correlation the CUBIC waveform with the POLY approximated waveform. In this particular example, the $\text{CUBIC}\circledast\text{POLY}$ curve lost $0.01~\text{dB}$ SNR. The $\text{CUBIC}\circledast\text{FFT}$ curve compares the CUBIC waveform with our FFT waveform method. For this example, the $\text{CUBIC}\circledast\text{FFT}$ compared similarly to the $\text{CUBIC}\circledast\text{POLY}$ curve as close as $0.001~\text{dB}$, indicating that our FFT method is on par with the accuracy of the Pseudo-Chebyshev polynomial method.
As can be seen in this example, the FFT approximation method and the Pseudo-Chebyshev method resulted in losses on the order of $0.01~\text{dB}$ SNR or less when compared to the cubic solution. The difference between each method resulted in peaks separated by $0.001~\text{dB}$. This analysis indicated that our FFT method is approximately as accurate as the Pseudo-Chebyshev polynomial approximation method for the parameters used to generate Fig. 2.
II-D Ionospheric Correction Implementation and Benchmarking
To demonstrate the computational rates of the methods considered, all tests in this paper were implemented in CUDA 12.3 and performed on the NVIDIA H100 GPU with a $3.25~\text{GHz}$ CPU processor connected via PCIe 5.0. This choice of hardware allowed us to parallelize the FFT, iFFT, and the phase shifting of each sample in our receive window. Figure 3 shows the benchmarking result of running (16) on a realistic waveform. Notably Fig. 3(b) shows benchmarks excluding the FFT and iFFT compute times.
((a))
((b))
Figure 3: Histograms showing the results from $10^{7}$ time trials where we performed the algorithm shown in (16). This test was run with $f_{0}=413~\text{MHz}$, $B=18~\text{MHz}$, a sample rate of $f_{s}=2.048~\text{GHz}$, $E=100e16~\text{electrons}/\text{m}^{2}$, $T=100~\mu\text{s}$, and had $2^{19}$ total I/Q samples. Figure 3(a) shows the rates when performing ionospheric correction during pulse compression, which averaged about $46~\mu\text{s}$ to complete and correlates to a rate of approximately $11.3~\text{Gigasamples}/\text{s}$. Figure 3(b) excludes the FFT and iFFT times and averages approximately $2200~\text{ns}$, which correlates to about $240~\text{Gigasamples}/\text{s}$.
As can be seen in Fig. 3, on our hardware we were able to complete the ionospheric correction considered in about $46~\mu\text{s}$; nearly $44~\mu\text{s}$ of which are from performing the required DFTs. We separate these test cases to demonstrate the efficiency that can be gained from bundling the ionospheric correction with other digital signal processing methods that require computations in the frequency domain. Common digital signal processing methods, including pulse compression, often require DFTs to be performed, and the ionospheric correction can be run in tandem with these other algorithms to allow for optimal signal processing rates. This efficiency makes our ionospheric correction method viable for use in digital processing in real time.
III Doppler Dispersion Correction
Similar to our approach to ionospheric distortion correction, we also investigated numeric Doppler dispersion correction methods on modern hardware. Because Doppler dispersion is already well-modeled analytically, this section will be focused on novel implementation of these numeric interpolation algorithms on the GPU. We first start with an overview of all relevant methods based on their time complexity and accuracy. We then provide an more-thorough analysis of the time performance and accuracy of one of the most-promising numerical methods that we present.
III-A Doppler Dispersion Time Dilation Methods